A laboratory technician needs to make a 45-liter batch of a 40% acid solution. How can the laboratory technician combine a batch of an acid solution that is pure acid with another that is 10% to get the desired concentration?
Answer by macston(5194)

.
P=amount of pure acid; T=amount of 10% solution
P+T=45L
T=45L-P
.
(1.00)P+(0.10)T=(0.40)45L
P+(0.10)(45L-P)=18L
P+4.5L-0.10P=18L
0.90P=13.5L
P=15L
ANSWER 1: He should use 15 liters of pure acid.
T=45L-P=45L-15L=30L
ANSWER 2: He should use 30 liters of 10% acid solution.
.
CHECK:
(1.00)P+(0.10)T=(0.40)45L
15L+(0.10)30L=18L
15L+3L=18L
18L=18L